ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 180, 2010

Bauer, F., D. Kraus, O. Roth and E. Wegert Beurling’s free boundary value problem in conformal geometry

223

Bell, J., D. Rogalski and S. J. Sierra The Dixmier–Moeglin equivalence for twisted homogeneous coordinate rings

461

Berarducci, A., M. Mamino and M. Otero Higher homotopy of groups definable in o-minimal structures

143

Berkovich, Y. Nonnormal and minimal nonabelian subgroups of a finite group

371

Chebotar, M. A., P.-H. Lee and E. R. Puczyłowski On Andrunakievich’s chain and Koethe’s problem

119

Chen, J.-F. and M.-L. Fang Normal families and shared functions of meromorphic functions

129

Chiang-Hsieh, H.-J., P.-F. Lee and H.-J. Wang The embedding of line graphs associated to the zero-divisor graphs of commutative rings

193

Choi, Y. S., H. J. Lee and H. G. Song Bishop’s theorem and differentiability of a subspace of C_{b}(K)

93

Cornelissen, G., F. Kato and A. Kontogeorgis The relation between rigid-analytic and algebraic deformation parameters for Artin–Schreier–Mumford curves

345

Costoya, C. Co-H-spaces and localization  69
Crestani, E. and P. Spiga Fixed-point-free elements in p-groups  413

Fang, M.-L. See J.-F. Chen Fernández-Alcober, G. A. and L. Legarreta The finite p-

groups with p conjugacy classes of non-normal subgroups

189

Gouëzel, S. Characterization of weak convergence of Birkhoff sums for Gibbs–Markov maps

1

Kandemir, M. and Ya. Yakubov Regular boundary value problems with a discontinuous coefficient, functional-multipoint conditions, and a linear spectral parameter

255

Kato, F. See G. Cornelissen Kontogeorgis, A. See G. Cornelissen Kraus, D. See F. Bauer Lee, H. J. See Y. S. Choi Lee, P.-F. See H.-J. Chiang-Hsieh Lee, P.-H. See M. A. Chebotar Legarreta, L. See G. A. Fernández-Alcober Liu, L. and D. Yang Characterizations of BMO associated with

Gauss measures via commutators of local fractional integrals

285

Mamino, M. See A. Berarducci Otero, M. See A. Berarducci

Pop, F. Pro-ℓ abelian-by-central Galois theory of prime divisors  43

Puczyłowski, E. R. See M. A. Chebotar

Pumplün, S. Classes of structurable algebras of skew-rank 1  425

Quéguiner-Mathieu, A. and J.-P. Tignol Algebras with involution that become hyperbolic over the function field of a conic

317

Rogalski, D. See J. Bell Rosenberg, M. Py-Calabi quasi-morphisms and quasi-states on ori-

entable surfaces of higher genus

163

Roth, O. See F. Bauer Shkarin, S. On the set of hypercyclic vectors for the differentiation

operator

271

Sierra, S. J. See J. Bell Song, H. G. See Y. S. Choi Spiga, P. See E. Crestani Tignol, J.-P See A. Quéguiner-Mathieu Wang, H.-J. See H.-J. Chiang-Hsieh Wegert, E. See F. Bauer Yakubov, Ya. See M. Kandemir Yang, D. See L. Liu